### v3

@@ 1 Introduction
$\begin{array}[]{l}\displaystyle\operatorname*{\textrm{argmin}\,}\limits_{\bm{W}\geq\bm{0},\bm{H}\geq\bm{0}}\bigg{\{}F(\bm{W},\bm{H})\coloneqq\dfrac{1}{n}\displaystyle\sum\limits_{j=1}^{n}\bigg{(}1-\dfrac{\langle\bm{m}_{:j},\bm{W}\bm{h}_{:j}\rangle}{\|\bm{m}_{:j}\|_{2}\|\bm{W}\bm{h}_{:j}\|_{2}}\bigg{)}\bigg{\}},\end{array}$ (Chordal-NMF)
[…]
Figure 1: Left: Picture of a rank-3 NMF. Data points (in black) that represent the columns of $\bm{M}$ encapsulated by a polyhedral cone generated by $r=3$ columns $\bm{w}_{1},\bm{w}_{2},\bm{w}_{3}$ (in red). Right: The sphere $\mathbb{S}^{2}_{+}$ in the nonnegative orthant +$\mathbb{R}^{3}_{+}$. The black dot is closer to the blue dot in Euclidean distance, but closer to the pink dot in chordal distance, which is equivalent to the geodesic arc length on the sphere.
@@
$\nabla f(\bm{x})=\big{(}\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\bm{A}^{\top}\bm{c}-\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})\big{)}\,\big{/}\,\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}.$
@@
2:     for $j=1,2,...,n$ (update $\bm{H}_{k+1}$ column-wise) do
3:         h-subproblem $\bm{h}_{:j,k+1}=\operatorname*{\textrm{argmax}\,}_{\bm{h}\geq\bm{0}}\big{\langle}\bm{m}_{:j},\bm{W}\bm{h}\big{\rangle}\big{/}\|\bm{W}\bm{h}\|_{2}.$
@@
5:     W-subproblem $\bm{W}_{k+1}=\operatorname*{\textrm{argmax}\,}_{\bm{W}\geq\bm{0}}\Bigg{\{}F(\bm{W};\bm{H})=\displaystyle\sum_{j=1}^{n}\dfrac{\langle\bm{m}_{:j},\bm{W}\bm{h}_{:j}\rangle}{\|\bm{W}\bm{h}_{:j}\|_{2}}\Bigg{\}}\Bigg{|}_{\bm{H}=\bm{H}_{k+1}}.$
@@ 1.1 Deriving and motivating chordal distance
  • Quotient space interpretation. From (2), we define a new objective in (Chordal-NMF) for measuring the distance between $\bm{m}_{:j}$ and $\bm{W}\bm{h}_{:j}$ by angle. The division of the norm $\|\bm{W}\bm{h}_{:j}\|_{2}$ in (Chordal-NMF) collapses the elements in the set $\{\bm{x}:\bm{x}=\bm{W}\bm{h}\}$ to a single point, leading to the chordal distance a purely angle-based distance by ignoring the length information.
@@
    $\bar{f}_{\text{chord}}:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow[0,2]\cup\{\infty\}\cup\{\pm\sqrt{-1}\infty\}\quad\text{as}\quad\bm{u},\bm{v}\mapsto\sqrt{2-2\langle\bm{u},\bm{v}\rangle/(\|\bm{u}\|_{2}\|\bm{v}\|_{2})},$
@@
    • $\bar{f}_{\text{chord}}$ is undefined at $\bm{u}=\bm{0}$ and/or $\bm{v}=\bm{0}$, it take $\infty$ or complex values ±$\pm\sqrt{-1}\infty$.
@@ 2 Background of optimization on manifold
[…]
@@ 2.2 Background of Riemannian optimization
[…]
@@
Retraction. A point $\bm{x}$ that originally sitting on a manifold $\mathcal{M}$ may go outside $\mathcal{M}$ after a gradient operation, so we need to pull it back onto $\mathcal{M}$. This can be achieved by a particular smooth map, known as the retraction $\mathcal{R}:T_{\bm{x}}\mathcal{M}\to\mathcal{M}$ with $\bm{v}\mapsto\mathcal{R}_{\bm{x}}(\bm{v})$. A retraction maps a point on the tangent space $T_{\bm{x}}\mathcal{M}$ onto $\mathcal{M}$ such that each curve $c(t)=\mathcal{R}_{x}(tv)$ satisfies $c(0)=\bm{x}$ and $c^{\prime}(0)=\bm{v}$ [18, Ch3, Def3.47]. There are different ways to perform retractions, such as exponential map and metric retractions [18, §3.6]. However, in general, to obtain the exponential map of a manifold, one requires to solve a differential equation. Thus, for computational efficiency, we consider the metric retraction $\mathcal{R}_{\bm{x}}(\bm{v})=\operatorname*{\textrm{argmin}\,}_{\bm{y}\in E}\|\bm{x}+\bm{v}-\bm{y}\|_{E}^{2}~{}\,\textrm{s.t.}\,~{}h(\bm{y})=0$ [18, §7.7]. Note that $\mathcal{R}_{\bm{x}}(\bm{v})$ is possibly non-unique and hard to compute. Among all the retractions we focus on
$\mathcal{R}_{\bm{x}}(\bm{v})\coloneqq(\bm{x}+\bm{v})/\|\bm{x}+\bm{v}\|_{E}.$ (Metric retraction)
@@ 2.3 Riemannian Multiplicative Update (RMU)
The element-wise operations as linear transformations in general do not preserve the tangent condition of a manifold into a point. However here we can still apply (Metric retraction) on $-\bm{\alpha}\odot\textrm{grad}f(\bm{x}_{k})$, the $\bm{x}_{k}$ terms got canceled and gives
@@
Proposition 2 also gives a convergence condition for $\textrm{grad}f(\bm{x}_{k})=\bm{0}$, which is equivalent to $\textrm{grad}^{-}f(\bm{x}_{k})=\textrm{grad}^{+}f(\bm{x}_{k})$. That is, we have a simple way to check the convergence of the sequence $\{\bm{x}_{k}\}_{k\in\mathbb{N}}$ by comparing $\textrm{grad}^{-}f(\bm{x}_{k}),\textrm{grad}^{+}f(\bm{x}_{k})$.
@@ 3.1 Column-wise h-subproblem over ellipsoid
$\operatorname*{\textrm{argmin}\,}_{\bm{h}}\big{\{}\phi(\bm{h})\coloneqq 1-\langle\bm{m}_{:j},\bm{W}\bm{h}\rangle\big{\}}\,\,\textrm{s.t.}\,\,\bm{h}\in\mathbb{R}^{r}_{+}\cap\mathcal{E}_{\bm{W}^{\top}\bm{W}}^{r-1},$ (h-manifold-subproblem)
@@
  • The set +$\mathbb{R}^{r}_{+}\cap\mathcal{E}_{\bm{W}^{\top}\bm{W}}^{r-1}$ is a nonsmooth manifold: +$\mathbb{R}^{r}_{+}$ has sharp corners at the boundary. See §2 for discussion on the issues caused by such non-smoothness.
@@
Retraction (Proposition 4) $\mathcal{R}_{\bm{\zeta}}(\bm{\xi})~{}=~{}(\bm{\zeta}+\bm{\xi})/\|\bm{\zeta}+\bm{\xi}\|_{\mathcal{E}}$
@@
Project $\bm{W}$ onto $T_{\bm{Z}}\mathcal{M}_{j}$ (Proposition 7) $\textrm{proj}_{T_{\bm{Z}}\mathcal{M}_{j}}(\bm{W})=\Big{(}\bm{I}-\dfrac{(\bm{Z}\bm{A}_{j})^{\otimes 2}}{\|\bm{Z}\bm{A}_{j}\|_{\textrm{F}}^{2}}\Big{)}\bm{W}$
Retraction (Proposition 8) $\mathcal{R}_{\bm{Z}}(\bm{W})=(\bm{Z}+\bm{W})\big{/}\|\bm{Z}+\bm{W}\|_{\bm{A}_{j}^{1/2}}$
@@
$\bm{\xi}-\dfrac{\langle\bm{A}\bm{\zeta},\bm{\xi}\rangle}{\|\bm{A}\bm{\zeta}\|_{2}^{2}}\bm{A}\bm{\zeta}~{}=~{}\Big{(}\bm{I}_{r}-\dfrac{(\bm{A}\bm{\zeta})^{\otimes 2}}{\|\bm{A}\bm{\zeta}\|_{2}^{2}}\Big{)}\bm{\xi}~{}\eqqcolon~{}\bm{P}\bm{\xi}.$
@@
$\mathcal{R}_{\bm{\zeta}}(\bm{\xi})~{}=~{}(\bm{\zeta}+\bm{\xi})\,/\,\|\bm{\zeta}+\bm{\xi}\|_{\mathcal{E}}\overset{T_{\bm{\zeta}}\mathcal{E}_{\bm{A}}^{r-1}}{~{}=~{}}(\bm{\zeta}+\bm{\xi})\,/\,\sqrt{1+\|\bm{\xi}\|_{\mathcal{E}}^{2}},$ ($\mathcal{R}$-ellipsoid)
@@
$\bm{h}_{:j,k+1}=\bm{z}/\|\bm{z}\|_{\mathcal{E}_{\bm{A}}^{r-1}},~{}\text{where}~{}\bm{z}=\bm{h}_{:j,k}\odot\textrm{grad}^{-}\phi(\bm{h}_{:j,k})[\bm{h}_{:j,k}]\oslash\textrm{grad}^{+}\phi(\bm{h}_{:j,k})[\bm{h}_{:j,k}].$
@@ 3.2 The finite sum matrix-wise W-subproblem
We do not solve (W-subproblem) with $n\geq 2$ using RMU (which is designed for $n=1$), we consider a consensus-averaging between all the manifolds that we call avgRMU. Furthermore, we try two alternative methods.
@@
$\nabla F(\bm{W})=\sum_{j=1}^{n}\Big{(}\bm{B}_{j}\,/\,\langle\bm{W},\bm{W}\bm{A}_{j}\rangle^{1/2}_{\textrm{F}}-\langle\bm{B}_{j},\bm{W}\rangle_{\textrm{F}}\bm{W}\bm{A}_{j}\,/\,\langle\bm{W},\bm{W}\bm{A}_{j}\rangle^{3/2}_{\textrm{F}}\Big{)}.$ (5)

### v4

@@ 1 Introduction
$\begin{array}[]{l}\displaystyle\hskip-8.5359pt\operatorname*{\textrm{argmin}\,}\limits_{\bm{W}\geq\bm{0},\bm{H}\geq\bm{0}}\Bigg\{F(\bm{W},\bm{H})\coloneqq\dfrac{1}{n}\displaystyle\sum\limits_{j=1}^{n}\bigg(1-\dfrac{\langle\bm{m}_{:j},\bm{W}\bm{h}_{:j}\rangle}{\|\bm{m}_{:j}\|_{2}\|\bm{W}\bm{h}_{:j}\|_{2}}\bigg)\Bigg\},\end{array}$ (Chordal-NMF)
where the objective function $F:\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}\to\mathbb{R}$ is defined as the chordal distance between $\bm{W}\bm{H}$ and $\bm{M}$, $\|\cdot\|_{2}$ is the Euclidean norm, $\langle\cdot,\cdot\rangle$ is the Euclidean inner product, and $\bm{m}_{:j}$ is the $j$th column of $\bm{M}$. The nonnegativity constraints $\geq\bm{0}$ is element-wise, where $\bm{0}$ is zero matrix of the appropriate size. For simplicity, we assume $\bm{M}\in\mathbb{R}^{m\times n}_{+}$ has no zero column, and $\bm{M}$ is normalized that $\|\bm{m}_{:j}\|_{2}=1$ in (Chordal-NMF). We give the motivation of using ray-to-ray distance in §1.1.
Figure 1: Left: Picture of a rank-3 NMF. Data points (in black) that represent the columns of $\bm{M}$ encapsulated by a polyhedral cone generated by $r=3$ columns $\bm{w}_{1},\bm{w}_{2},\bm{w}_{3}$ (in red). Right: The sphere $\mathbb{S}^{2}_{+}$ in the nonnegative orthant $\mathbb{R}^{3}_{+}$. The black dot is closer to the blue dot in Euclidean distance, but closer to the pink dot in chordal distance, which is equivalent to the geodesic arc length on the sphere.
@@
$\nabla f(\bm{x})=\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\bm{A}^{\top}\bm{c}-\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})}{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}.$
@@
2:  for j=1,2,,nj=1,2,...,n (update $\bm{H}_{k+1}$ column-wise) do
3:   h-subproblem $\bm{h}_{:j,k+1}=\operatorname*{\textrm{argmax}\,}_{\bm{h}\geq\bm{0}}\dfrac{\langle\bm{m}_{:j},\bm{W}\bm{h}\big\rangle}{\|\bm{W}\bm{h}\|_{2}}.$
@@
5:  W-subproblem $\bm{W}_{k+1}=\operatorname*{\textrm{argmax}\,}_{\bm{W}\geq\bm{0}}\hskip-2.84526pt\Bigg\{F(\bm{W};\bm{H})=\displaystyle\sum_{j=1}^{n}\dfrac{\langle\bm{m}_{:j},\bm{W}\bm{h}_{:j}\rangle}{\|\bm{W}\bm{h}_{:j}\|_{2}}\Bigg\}\Bigg|_{\bm{H}=\bm{H}_{k+1}}.$
@@ 1.1 Deriving and motivating chordal distance
  • Quotient space interpretation. From (2), we define a new objective in (Chordal-NMF) for measuring the distance between $\bm{m}_{:j}$ and $\bm{W}\bm{h}_{:j}$ by angle. The division of the norm $\|\bm{W}\bm{h}_{:j}\|_{2}$ in (Chordal-NMF) collapses all vectors $\bm{x}$ that are proportional to $\bm{W}\bm{h}$ onto a single point on the unit sphere. In this way, the chordal distance becomes purely angle-based, since the length information is ignored.
@@
    $\bar{f}_{\text{chord}}:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow[0,2]\cup\{\infty\}\cup\{\pm\sqrt{-1}\infty\}\quad\text{as}\quad\bm{u},\bm{v}\mapsto\sqrt{2-2\dfrac{\langle\bm{u},\bm{v}\rangle}{\|\bm{u}\|_{2}\|\bm{v}\|_{2}}},$
@@
    • $\bar{f}_{\text{chord}}$ is undefined at $\bm{u}=\bm{0}$ and/or $\bm{v}=\bm{0}$, it take $\infty$ or complex values $\pm\sqrt{-1}\infty$.
@@
Given the above discussion and the fact that the chordal distance is non-Euclidean, we employ Riemannian optimization methods in this work.
@@ 2 Background of optimization on manifold
$\mathcal{P}_{0}:\operatorname*{\textrm{argmin}\,}_{\bm{x}}f(\bm{x})~\,\textrm{s.t.}\,~\bm{x}\in\mathcal{M},\,\bm{x}\in\mathcal{X}~~\overset{\text{restriction}}{\iff}~~\mathcal{P}_{1}:\operatorname*{\textrm{argmin}\,}_{\bm{x}\in\mathcal{M}}f(\bm{x})~\,\textrm{s.t.}\,~\bm{x}\in\mathcal{X},$
@@ 2.2 Background of Riemannian optimization
Submanifold and ambient space. Let $E$ be an Euclidean space with $dim(E)>k$, with inner product $\langle\cdot,\cdot\rangle$, and induced norm $\|\cdot\|$. Considering the smooth function $h:E\to\mathbb{R}^{k}$, with $\textrm{D}h(\bm{x})[\bm{v}]=\langle\textrm{grad}h(\bm{x}),\bm{v}\rangle_{E}$ its differential with full rank $k$ for all $\bm{x}$ such that $h(\bm{x})=\bm{0}_{k}$, the set $\mathcal{M}\coloneqq\{\bm{x}\in E~|~h(\bm{x})=\bm{0}_{k}\}$ is an embedded submanifold of $E$ of dimension $\dim E-k$ (see [18, Ch7.7]). In this work we focus on $h$ in the form $h(\bm{x})=\langle\bm{x},\bm{x}\rangle_{E}-1$. For a manifold $\mathcal{M}$ in $E$, we call $E$ the ambient space of $\mathcal{M}$. In this work, all ambient spaces are some specific linear vector space $E$, such as $\mathbb{R}^{r}$ and $\mathbb{R}^{m\times r}$.
@@
Retraction. A point $\bm{x}$ that is originally sitting on a manifold $\mathcal{M}$ may go outside $\mathcal{M}$ after a gradient operation, so we need to pull it back onto $\mathcal{M}$. This can be achieved by a particular smooth map, known as the retraction $\mathcal{R}:T_{\bm{x}}\mathcal{M}\to\mathcal{M}$ with $\bm{v}\mapsto\mathcal{R}_{\bm{x}}(\bm{v})$. A retraction maps a point on the tangent space $T_{\bm{x}}\mathcal{M}$ onto $\mathcal{M}$ such that each curve $c(t)=\mathcal{R}_{x}(tv)$ satisfies $c(0)=\bm{x}$ and $c^{\prime}(0)=\bm{v}$ [18, Ch3, Def3.47]. There are different ways to perform retractions, such as exponential map and metric retractions [18, §3.6]. However, in general, to obtain the exponential map of a manifold, one requires to solve a differential equation. Thus, for computational efficiency, we consider the metric retraction $\mathcal{R}_{\bm{x}}(\bm{v})=\operatorname*{\textrm{argmin}\,}_{\bm{y}\in E}\|\bm{x}+\bm{v}-\bm{y}\|_{E}^{2}~\,\textrm{s.t.}\,~h(\bm{y})=0$ [18, §7.7]. Note that $\mathcal{R}_{\bm{x}}(\bm{v})$ is possibly non-unique and hard to compute. Among all the retractions we focus on
$\mathcal{R}_{\bm{x}}(\bm{v})\coloneqq\dfrac{\bm{x}+\bm{v}}{\|\bm{x}+\bm{v}\|_{E}}.$ (Metric retraction)
@@ 2.3 Riemannian Multiplicative Update (RMU)
The element-wise operations as linear transformations in general do not preserve the tangent condition of a manifold into a point. However here we can still apply (Metric retraction) on $\bm{\alpha}\odot\textrm{grad}f(\bm{x}_{k})$, the $\bm{x}_{k}$ terms got canceled and gives
@@
Proposition 2 also gives a convergence condition to a critical point:
  • If the solution lies strictly inside the nonnegative orthant, this corresponds to $\textrm{grad}f(\bm{x}_{k})=\bm{0}$, which is equivalent to $\textrm{grad}^{-}f(\bm{x}_{k})=\textrm{grad}^{+}f(\bm{x}_{k})$. That is, the convergence of the sequence $\{\bm{x}_{k}\}_{k\in\mathbb{N}}$ can be easily checked by comparing $\textrm{grad}^{-}f(\bm{x}_{k})$ and $\textrm{grad}^{+}f(\bm{x}_{k})$.
  • Otherwise, in the constrained setting, convergence is characterized by the KKT conditions on $\mathcal{M}\cap\mathbb{R}^{n}_{+}$.
@@ 3.1 Column-wise h-subproblem over ellipsoid
$\operatorname*{\textrm{argmin}\,}_{\bm{h}}\Big\{\phi(\bm{h})\coloneqq 1-\langle\bm{m}_{:j},\bm{W}\bm{h}\rangle\Big\}~\,\textrm{s.t.}\,~\bm{h}\in\mathbb{R}^{r}_{+}\cap\mathcal{E}_{\bm{W}^{\top}\bm{W}}^{r-1},$ (h-manifold-subproblem)
@@
  • The set $\mathbb{R}^{r}_{+}\cap\mathcal{E}_{\bm{W}^{\top}\bm{W}}^{r-1}$ is a nonsmooth manifold: $\mathbb{R}^{r}_{+}$ has sharp corners at the boundary. See §2 for discussion on the issues caused by such non-smoothness.
@@
Retraction (Proposition 4) $\mathcal{R}_{\bm{\zeta}}(\bm{\xi})~=~\dfrac{\bm{\zeta}+\bm{\xi}}{\|\bm{\zeta}+\bm{\xi}\|_{\mathcal{E}}}$
@@
Project $\bm{W}$ onto $T_{\bm{Z}}\mathcal{M}_{j}$ (Proposition 7) $\textrm{proj}_{T_{\bm{Z}}\mathcal{M}_{j}}(\bm{W})=\bigg(\bm{I}-\dfrac{(\bm{Z}\bm{A}_{j})^{\otimes 2}}{\|\bm{Z}\bm{A}_{j}\|_{\textrm{F}}^{2}}\bigg)\bm{W}$
Retraction (Proposition 8) $\mathcal{R}_{\bm{Z}}(\bm{W})=\dfrac{\bm{Z}+\bm{W}}{\|\bm{Z}+\bm{W}\|_{\bm{A}_{j}^{1/2}}}$
@@
$\bm{\xi}-\dfrac{\langle\bm{A}\bm{\zeta},\bm{\xi}\rangle}{\|\bm{A}\bm{\zeta}\|_{2}^{2}}\bm{A}\bm{\zeta}~=~\bigg(\bm{I}_{r}-\dfrac{(\bm{A}\bm{\zeta})^{\otimes 2}}{\|\bm{A}\bm{\zeta}\|_{2}^{2}}\bigg)\bm{\xi}~\eqqcolon~\bm{P}\bm{\xi}.$
@@
$\mathcal{R}_{\bm{\zeta}}(\bm{\xi})~=~\dfrac{\bm{\zeta}+\bm{\xi}}{\|\bm{\zeta}+\bm{\xi}\|_{\mathcal{E}}}\overset{T_{\bm{\zeta}}\mathcal{E}_{\bm{A}}^{r-1}}{~=~}\dfrac{\bm{\zeta}+\bm{\xi}}{\sqrt{1+\|\bm{\xi}\|_{\mathcal{E}}^{2}}},$ ($\mathcal{R}$-ellipsoid)
@@
$\bm{h}_{:j,k+1}=\dfrac{\bm{z}}{\|\bm{z}\|_{\mathcal{E}_{\bm{A}}^{r-1}}},~\text{where}~\bm{z}=\bm{h}_{:j,k}\odot\textrm{grad}^{-}\phi(\bm{h}_{:j,k})[\bm{h}_{:j,k}]\oslash\textrm{grad}^{+}\phi(\bm{h}_{:j,k})[\bm{h}_{:j,k}].$
@@ 3.2 The finite sum matrix-wise W-subproblem
We do not solve (W-subproblem) with $n\geq 2$ using RMU (which is designed for $n=1$), For illustration, we also considered a consensus-averaging approach between all the manifolds, which we call avgRMU. The motivation was that the manifolds can conflict with each other in RMU, so averaging provides a natural “relaxation” to fit RMU in the intersection of the manifolds. We remark that later we found that experiments show that this approach performs worse than the Euclidean method, and we do not use it as our main solution.
Furthermore, we try two alternative methods.
@@
$\nabla F(\bm{W})~=~\sum_{j=1}^{n}\dfrac{\bm{B}_{j}}{\langle\bm{W},\bm{W}\bm{A}_{j}\rangle^{1/2}_{\textrm{F}}}-\dfrac{\langle\bm{B}_{j},\bm{W}\rangle_{\textrm{F}}\bm{W}\bm{A}_{j}}{\langle\bm{W},\bm{W}\bm{A}_{j}\rangle^{3/2}_{\textrm{F}}}.$ (5)